# Efficient Large-Scale Urban Parking Prediction: Graph Coarsening Based on Real-Time Parking Service Capability

### v1

@@ III-C Parking graph Coarsening
0:  $A_{c}\in\mathbb{R}^{N_{c}\times N_{c}},\text{index}\leftarrow\left[\ \right]$
1:  $N_{c}\leftarrow\text{CoarsenedNodesNum}(n,N)$
2:  $\mathcal{L}\leftarrow\text{NormalizedLaplacian}(A_{\text{combined}})$
3:  $\lambda,u\leftarrow\text{EigenValueAndVector}(L,N_{c})$
4:  for $i\leftarrow 1\text{ to }N$ do
5:     for $j\leftarrow 1\text{ to Neighbor}(i)$ do
6:        $A_{c}^{\text{temp}},\text{index}\leftarrow\text{ParkingSubZoneClustering}(u_{i},u_{j})$
7:     end for
8:  end for
9:  for $i\leftarrow 1\text{ to }N$ do
10:     for $j\leftarrow 1\text{ to }N$ do
11:        $A_{c}\leftarrow\arg\min\text{SD}(A_{\text{combined}}(i,j),A_{c}(i,j))$
@@
14:  return $A_{c},\text{index}$
III-D Prediction framework based on coarsened parking graphs
@@
Figure 2: The structure of Urban Parking Prediction Framework.
@@
The procedure is detailed as follows: As depicted in Fig. 2, firstly, the parking data $X^{\text{low}}\in\mathbb{R}^{T\times N\times F^{\text{low}}}$ undergoes encoding with the residual blocks within the encoder[44], as indicated in Formula 15. Owing to the reduction in the number of features as the quantity of residual blocks increases, it allows the feature data $X_{c}\in\mathbb{R}^{T\times N_{c}\times F^{\text{low}}}$ corresponding to the coarsened graph to be represented within a low-dimensional feature space.
$X_{c}(i)=\left(X^{\text{low }}*f\right)(i)=\sum_{j=0}^{k-1}f(j)X^{\text{low }}(i-j),$ (15)
@@
Next, we input the adjacency matrix $A_{c}$ of coarsened graph and feature matrix $X_{c}$ into a spatio-temporal graph convolutional model. This paper takes T-GCN[33] as an example, which captures the spatio-temporal dependencies of the parking coarsened graph through 2 layers of GCN and 1 layer of GRU, finally obtaining the coarsened graph’s predicted results $Y_{c}\in\mathbb{R}^{T\times N_{c}\times F^{\text{low}}}$ through a linear transformation, as seen in Formulas 16, 17, 18, 19, and 20, where $W$ and $b$ represent the weights and biases of T-GCN, respectively, and $u_{t}$, $r_{t}$, $c_{t}$, $h_{t}$ denote the reset gate, update gate, candidate hidden state, and current hidden state at time $t$,respectively. $gc\left(\cdot\right)$ denotes the graph convolution operation, and tanh is the activation function.
$u_{t}=\sigma\left(W_{u}\cdot\left[gc\left(X_{c},A_{c}\right),h_{t-1}\right]+b_{u}\right),$ (16)
$r_{t}=\sigma\left(W_{r}\cdot\left[gc\left(X_{c},A_{c}\right),h_{t-1}\right]+b_{r}\right),$ (17)
$c_{t}=\tanh\left(W_{c}\cdot\left[gc\left(X_{c},A_{c}\right),\left(r_{t},h_{t-1}\right)\right]+b_{c}\right),$ (18)
@@ IV-A Experiment Setup
In the experimental parameter design, we use data from the past 12 time steps to predict data for the next 1, 2, and 4 steps, which corresponds to using the past 3 hours of parking traffic to predict the parking traffic for the next 15 minutes, 30 minutes, and 60 minutes, respectively. Additionally, to evaluate the predictive performance, we selected three widely used baseline models in the parking prediction field for comparison. Additionally, to explore the effectiveness of graph coarsening techniques, we also employed the recently popular graph sparsification methods as a benchmark:
  • T-GCN[33]: A classical model for traffic prediction that combines Graph Convolutional Networks (GCNs) and Gated Recurrent Units (GRU) to establish spatio-temporal correlations among traffic data.
  • STGCN[46]: Utilizes multiple ST-Conv blocks to model multi-scale traffic graphs, proven to effectively capture comprehensive spatio-temporal correlations.
  • STSGCN[47]: By stacking multiple STSGCL blocks for synchronous spatio-temporal modeling, it can effectively capture complex local spatio-temporal traffic correlations.
@@
For the training of various models, we utilized V100 graphics cards and implemented early stopping and dynamic learning rate adjustment mechanisms to optimize the training process. To ensure completeness and transparency of information, we have listed the configuration details of the models along with all training-related hyperparameters in Table \@slowromancapii@.
@@
STSGCN GCNs per Module Spatio-temporal GCNs Layers(STSGCL) 64 1e-5 Adam Huber Loss 1e-4 200
@@ IV-C Selection of Coarsening Ratio
As seen in Fig. 6, as the coarsening ratio decreases from 1 to 0.6, the error growth rate of the model predictions remains relatively stable. This indicates that within this range of coarsening ratio, the model primarily loses redundant and secondary parking information, the absence of which has limited impact on overall prediction accuracy. However, as the coarsening ratio is further reduced, the prediction error of the models rapidly increases, particularly when the coarsening ratio is adjusted to 0.2, introducing calculation errors of 77.2%, 64.2%, and 61.8%, respectively. This suggests that crucial parking information (such as the highly volatile characteristics of parking lots) has been stripped away in the coarsening process.
@@
It’s noteworthy that the curve of model training duration presents a different trend. With the reduction of the coarsening ratio, the initial training time decreases rapidly, showing that coarsening can significantly reduce the computational burden of the model. However, after reaching a coarsening ratio of 0.6, the trend of reduction in training duration becomes more gradual, indicating that further coarsening does not significantly enhance efficiency. Specifically, in the coarsened graph, nodes of parking lots with significantly different characteristics may be merged, leading to a loss of crucial detail information that was originally observable. Consequently, the model’s ability to distinguish these differences is impaired, thereby affecting its predictive and generalization capabilities.
@@ IV-E Scalability Results
To verify whether our proposed framework can adapt to parking graphs of different scales, we simulate the expansion of the parking graph scale by gradually increasing the number of parking lots. In the experiments, the coarsening ratio is set to 0.6, with the prediction of parking occupancy rates for the next 15 minutes during the evening rush hour as the evaluation task. This approach allows us to observe changes in the model’s prediction accuracy and convergence speed on parking graphs of various sizes. Synthesizing Fig. 7(a) and 7(b), it can be observed that when the scale of the parking graph is small, our framework shows relatively conservative improvements in terms of accuracy and efficiency, with an average accuracy improvement of about 11.1% and a training efficiency improvement of about 23.7%. Importantly, as the scale of the graph expands, the performance advantages of our prediction framework become particularly pronounced. Unlike the baseline models, which exhibit a higher error growth rate and exponential increase in iteration costs, our framework demonstrates a stable growth trend, significantly surpassing traditional methods in terms of error control and training efficiency.
@@
IV-F Ablation Results
@@
Fig. 8(a) shows the experimental results of parking predictions using various types of adjacency matrices. Clearly, among all baseline methods, the ParkingRank attention matrix constructed by PRGAT exhibits the best performance, achieving an average reduction in relative error of about 38.3%. PRGAT, by integrating the attention coefficient matrix with the parking spatio-temporal transition matrix, delves deep into the feature and spatial dependencies between parking lots, effectively restoring real parking demand scenarios. In contrast, traditional static adjacency matrices, built solely on spatial location information, overlook the intrinsic characteristics of parking lots and their interactions with the surrounding environment. Often, they fail to fully capture the complex interactions and dynamic changes between parking lots, limiting the authenticity and dynamism of the scenarios captured by the model, and thus performing poorly in downstream prediction tasks. Moreover, compared to the traditional GAT mechanism, another advantage of PRGAT is that it not only offers a more detailed and dynamic method to represent parking graphs, but it also intuitively reflects which parking lots have stronger comprehensive service capabilities at specific moments, thus becoming a key basis for making parking decisions.
@@
Figure 8: Ablation Study on Future 15-Minute Parking Prediction (a) Graph Construction (b) Autoencoder.

### v2

@@ III-D Parking Graph Coarsening
0:  $A_{c}\in\mathbb{R}^{N_{c}\times N_{c}},\text{index}$
[…]
2:  $N_{c}\leftarrow\text{CoarsenedNodesNum}(n,N)$
3:  $\mathcal{L}\leftarrow\text{NormalizedLaplacian}(A_{\text{combined}})$
4:  $\lambda,u\leftarrow\text{EigenValueAndVector}(L,N_{c})$
5:  for $i\leftarrow 1\text{ to }N$ do
6:     for $j\leftarrow 1\text{ to Neighbor }(i)$ do
7:        $d_{c}(i,j)=\frac{\sum_{i=1,j=1}^{n}u_{i}u_{j}}{\sqrt{\sum_{i=1}^{n}\left(u_{i}\right)^{2}}\sqrt{\sum_{j=1}^{n}\left(u_{j}\right)^{2}}}$
[…]
9:        for $q\leftarrow 1\text{ to Neighbor }(temp)$ do
[…]
11:        end for
@@
14:  for $i\leftarrow 1\text{ to }N$ do
15:     for $j\leftarrow 1\text{ to }N$ do
16:        $A_{c}\leftarrow\arg\min\text{SD}(A_{\text{combined}}(i,j),A_{c}(i,j))$
17:     end for
18:  end for
19:  return $A_{c},\text{index}$
III-E Prediction Framework Based On Coarsened Parking Graphs
@@
Figure 2: The structure of TCN-AE.
@@
The procedure is detailed as follows: As depicted in Fig. 2, firstly, the parking data $X^{\text{low}}\in\mathbb{R}^{T\times N\times F^{\text{low}}}$ undergoes encoding with the residual blocks within the encoder[40], as indicated in Formula 15. Owing to the reduction in the number of features as the quantity of residual blocks increases, it allows the feature data $X^{c}\in\mathbb{R}^{T\times N_{c}\times F^{\text{low}}}$ corresponding to the coarsened graph to be represented within a low-dimensional feature space.
$X^{\text{c}}(i)=\left(X^{\text{low }}*f\right)(i)=\sum_{j=0}^{k-1}f(j)X^{\text{low }}(i-j),$ (15)
@@
Figure 3: The structure of T-GCN.
Next, we input the adjacency matrix $A_{c}$ of coarsened graph and feature matrix $X^{c}$ into a spatiotemporal graph convolutional model. This paper takes T-GCN[30] as an example, as shown in Fig. 3, which captures the spatiotemporal dependencies of the parking coarsened graph through 2 layers of GCN and 1 layer of GRU, finally obtaining the coarsened graph’s predicted results $Y_{c}\in\mathbb{R}^{T\times N_{c}\times F^{\text{low}}}$ through a linear transformation, as seen in Formulas 16, 17, 18, 19, and 20, where $W$ and $b$ represent the weights and biases of T-GCN, respectively, and $u_{t}$, $r_{t}$, $c_{t}$, $h_{t}$ denote the reset gate, update gate, candidate hidden state, and current hidden state at time $t$,respectively. $gc\left(\cdot\right)$ denotes the graph convolution operation, and tanh is the activation function.
$u_{t}=\sigma\left(W_{u}\cdot\left[gc\left(X^{c},A_{c}\right),h_{t-1}\right]+b_{u}\right),$ (16)
$r_{t}=\sigma\left(W_{r}\cdot\left[gc\left(X^{c},A_{c}\right),h_{t-1}\right]+b_{r}\right),$ (17)
$c_{t}=\tanh\left(W_{c}\cdot\left[gc\left(X^{c},A_{c}\right),\left(r_{t},h_{t-1}\right)\right]+b_{c}\right),$ (18)
@@
The overall parking prediction framework of this paper is shown in Fig. 4.
Figure 4: The structure of Urban Parking Prediction Framework.
@@ IV-A Experiment Setup
In the experimental parameter design, we use data from the past 12 time steps to predict data for the next 1, 2, and 4 steps, which corresponds to using the past 3 hours of parking traffic to predict the parking traffic for the next 15 minutes, 30 minutes, and 60 minutes, respectively. Additionally, to evaluate the performance of our proposed urban parking prediction framework, we selected several widely-used baseline models in the field of parking prediction, as well as some recent popular graph reduction methods for comparison:
  • T-GCN[32]: A classical model for traffic flow prediction that combines Graph Convolutional Networks (GCNs) and Gated Recurrent Units (GRU) to capture the spatiotemporal correlations in traffic data, effectively modeling both spatial dependencies between traffic nodes and temporal patterns within the traffic time series.
  • STGCN[44]: Utilizes multiple ST-Conv blocks that integrate temporal and spatial convolutions to model multi-scale traffic networks, proven to effectively capture comprehensive spatiotemporal correlations.
  • STSGCN[45]: By stacking multiple STSGCL blocks for synchronous spatiotemporal modeling, it can effectively capture complex local spatiotemporal traffic correlations.
@@
For the training of different models, we used a V100 GPU for development and incorporated early stopping and dynamic learning rate adjustment strategies to optimize training performance. To ensure completeness and transparency, all model configurations and training-related hyperparameter settings are detailed in Table \@slowromancapii@. In these configurations, the structural parameters of each model were kept at their default settings, while other hyperparameters, such as learning rate and weight decay, were selected using grid search.
@@
STSGCN GCNs per Module Spatiotemporal GCNs Layers(STSGCL) 64 1e-5 Adam Huber Loss 1e-4 200
@@ IV-C Selection of Coarsening Ratio
As shown in Fig. 8, when the coarsening ratio decreases from 1 to 0.6, the error growth rate of the model predictions remains relatively stable. This indicates that within this range, key features are effectively preserved, and the model primarily loses redundant and secondary parking information, whose absence has limited impact on overall prediction accuracy. However, as the coarsening ratio is further reduced, prediction error rises sharply. For instance, at a coarsening ratio of 0.2, calculation errors of 77.2%, 64.2%, and 61.8% are introduced across different model configurations. This suggests that crucial parking information, including the highly variable characteristics of certain parking lots, has been stripped away during the coarsening process, which significantly degrades model performance.
@@
It’s noteworthy that the curve of model training duration presents a different trend. Initially, as the coarsening ratio reduces, training time decreases rapidly, demonstrating that coarsening can effectively alleviate the model’s computational burden. However, once the coarsening ratio reaches 0.6, the reduction in training time becomes more gradual, suggesting that further coarsening does not yield significant efficiency gains. This shift indicates that the coarsening process has started merging nodes of parking lots with distinct characteristics, resulting in the loss of critical detail information that was previously observable. Under these conditions, the model’s ability to differentiate between variations in parking lot service capacities gradually deteriorates, making it less efficient in learning the complex features of urban parking networks in a lower-information setting. This, in turn, impacts its prediction accuracy and generalization performance.
@@ IV-E Scalability Results
To verify whether our proposed framework can adapt to parking graphs of different scales, we simulate the expansion of the parking graph scale by gradually increasing the number of parking lots. In the experiments, the coarsening ratio is set to 0.6, with the prediction of parking occupancy rates for the next 15 minutes during the evening rush hour as the evaluation task. This approach allows us to observe changes in the model’s prediction accuracy and convergence speed on parking graphs of various sizes. Synthesizing Fig. 9(a) and 9(b), it can be observed that when the scale of the parking graph is small, our framework shows relatively conservative improvements in terms of accuracy and efficiency, with an average accuracy improvement of about 11.1% and a training efficiency improvement of about 23.7%. Importantly, as the scale of the graph expands, the performance advantages of our prediction framework become particularly pronounced. Unlike the baseline models, which exhibit a higher error growth rate and exponential increase in iteration costs, our framework demonstrates a stable growth trend, significantly surpassing traditional methods in terms of error control and training efficiency.
@@
IV-F Robustness Analysis
To validate the robustness of our framework under different parking patterns and regional layouts, we further refined the experimental dataset and conducted robustness analysis on the subdivided datasets. Specifically, using the parking data of Bao’an District as the base, we divided the dataset into three smaller subsets based on differences in parking lot service coverage. These subsets correspond to regions primarily dominated by shopping centers, office complexes, and areas with a relatively balanced distribution of parking lot types. Each subset contains approximately 200 parking lots, as detailed in Table \@slowromancapiv@. In the experimental setup, we set the prediction step to 1, the coarsening rate to 0.6, and used MAE and MAPE as the main evaluation metrics.
TABLE IV: Dataset Settings for Different Parking Types
Dataset Type
Shopping
Center
Parking Lots
Office
Complex
Parking Lots
Residential
Compound
Parking Lots
Commercial Dataset 133 30 37
Office Dataset 34 122 44
Balanced Dataset 72 65 63
As shown in Table \@slowromancapv@, the prediction results for different types of parking datasets demonstrate that our framework achieves higher prediction accuracy on datasets with distinct regional characteristics compared to parking networks with evenly distributed parking lot types. Specifically, in parking networks dominated by commercial area parking lots, our framework improves prediction accuracy by 11.59% in terms of MAE. Similarly, for parking networks dominated by office area parking lots, the MAE prediction error is reduced by an average of 13.18%. In terms of MAPE, our framework achieves an average improvement of 15.06% in prediction performance on these two types of parking networks with highly concentrated service characteristics.
TABLE V: Comparison of 15-Minute Ahead Prediction Performance Across Different Types of Parking Datasets
Datasets
PRGAT+
Coarsening+
TGCN+TCN-AE
PRGAT+
Coarsening+
STGCN+TCN-AE
PRGAT+
Coarsening+
STSGCN+TCN-AE
MAE MAPE(%) MAE MAPE(%) MAE MAPE(%)
Commercial
Dataset
3.7208 6.14 2.2168 5.58 4.0399 5.63
Office
Dataset
3.5564 6.37 2.1077 5.42 4.1988 5.80
Balanced
Dataset
4.3507 7.44 2.4326 6.71 4.5624 6.44
The performance improvement is primarily attributed to the fact that parking networks with distinct regional characteristics often exhibit clearer parking demand and behavior patterns. For example, the demand fluctuations for commercial area parking lots are mainly concentrated on weekends, while the peak demand for office area parking lots typically occurs during morning and evening rush hours. Such parking networks with well-defined overall trends provide a solid foundation for the optimization of the PRGAT module in our framework. Specifically, the characteristics of these networks enable the PRGAT module to accurately identify and enhance the feature representation of critical nodes while weakening the influence of secondary and redundant nodes on model performance. This allows the construction of an adjacency matrix that better reflects actual parking behavior patterns. The adjacency matrix, in turn, provides the downstream spatiotemporal graph convolutional model with clearer learning objectives, thereby improving the model’s prediction performance.
In contrast, for parking networks with evenly distributed types and less pronounced demand characteristics, the correlations between nodes are relatively uniform, which limits the effectiveness of the PRGAT module in constructing the adjacency matrix and results in a smaller improvement in the framework’s predictive performance.Through these experimental validations, our framework is better suited for parking networks with distinct regional characteristics, whereas the overall performance improvement is relatively limited for networks with more evenly distributed parking types.
IV-G Ablation Results
@@
Fig. 10(a) shows the experimental results of parking predictions using various types of adjacency matrices. Clearly, among all baseline methods, the ParkingRank attention matrix constructed by PRGAT exhibits the best performance, achieving an average reduction in relative error of about 38.3%. PRGAT, by integrating the attention coefficient matrix with the parking spatiotemporal transition matrix, delves deep into the feature and spatial dependencies between parking lots, effectively restoring real parking demand scenarios. In contrast, traditional static adjacency matrices, built solely on spatial location information, overlook the intrinsic characteristics of parking lots and their interactions with the surrounding environment. Often, they fail to fully capture the complex interactions and dynamic changes between parking lots, limiting the authenticity and dynamism of the scenarios captured by the model, and thus performing poorly in downstream prediction tasks. Moreover, compared to the traditional GAT mechanism, another advantage of PRGAT is that it not only offers a more detailed and dynamic method to represent parking graphs, but it also intuitively reflects which parking lots have stronger comprehensive service capabilities at specific moments, thus becoming a key basis for making parking decisions.
@@
Figure 10: Ablation Study on Future 15-Minute Parking Prediction. (a) Graph Construction (b) Autoencoder.
@@
IV-H Node Error Analysis
To further validate the contribution of the ParkingRank Graph Attention Module to the accuracy of parking prediction tasks, we conducted a node error comparison experiment. Specifically, we used the original adjacency matrix and the attention matrix generated by the pre-trained ParkingRank Graph Attention Module as inputs to the spatiotemporal graph convolution model and recorded the average prediction error of each parking lot node during training to evaluate the impact of different adjacency matrices on model performance. The experiment utilized data from 100 parking lots in Bao’an District, with T-GCN as the parking prediction model and a prediction step set to 1.
Figure 11: Heatmap Visualization. (a) Original Adjacency Matrix (b) Node Average Prediction Error Distribution under the Original Adjacency Matrix (c) ParkingRank Attention Matrix (d) Node Average Prediction Error Distribution under the ParkingRank Attention Matrix.
Fig. 11(a) shows the heatmap visualization of the original adjacency matrix, where each cell represents the spatial relationships between different parking lots. From the figure, it can be observed that the weight distribution is relatively uniform, lacking evident traffic patterns and presenting a certain level of noise. Fig. 11(b) illustrates the prediction error distribution of different parking lot nodes during the training process when the original adjacency matrix is used as input for the spatiotemporal graph convolution model. It is evident that the prediction errors for all parking lot nodes are generally high, indicating that the model struggles to effectively capture the spatiotemporal relationships between parking lots. For certain important parking lots in the area (e.g., those near densely packed commercial zones or office buildings, highlighted in deep red in the figure), the prediction errors are particularly pronounced. This phenomenon further demonstrates that the original adjacency matrix fails to effectively reflect the underlying traffic patterns and spatial features of the parking network, thereby limiting the model’s prediction performance.
Fig. 11(c) presents the heatmap visualization of the ParkingRank Attention Adjacency Matrix constructed by the ParkingRank Graph Attention Module. Compared to Fig. 11(a), the weight distribution in Fig. 11(c) is more heterogeneous and structured. Specifically, the distribution characteristics of important parking lots are more evident in Fig. 11(c), as reflected by larger weight accumulations in the corresponding rows or columns, with overall darker colors. This feature indicates that certain parking lots are likely located in commercial centers or other areas with high parking demand, showing stronger associations with nearby parking lots. As a result, distinct cluster-like structures are formed in the corresponding rows or columns of the adjacency matrix. These cluster structures reflect the local correlations between parking lots, effectively revealing the actual characteristics of the parking network and its spatiotemporal interaction patterns.
Fig. 11(d) illustrates the prediction error distribution of different parking lot nodes during the training process when the ParkingRank Attention Matrix is used as input. Compared to Fig. 11(b), where the original adjacency matrix is used, the error levels of all parking lot nodes are significantly reduced. In particular, for nodes at the center of cluster structures (i.e., parking lot nodes with darker colors in Fig. 11(b)), the model effectively reduces the average prediction error by 45.1% during each training epoch. This result demonstrates that the ParkingRank Attention Matrix enables the spatiotemporal graph convolution model to achieve higher learning efficiency and lower error levels in regions with distinct features. Specifically, the ParkingRank Attention Matrix optimizes the weight distribution of the original adjacency matrix, more accurately reflecting the relationships between important parking lots and their surrounding lots. This refined weight allocation not only strengthens the relational representation of key nodes but also reduces the influence of minor and redundant nodes, allowing the model to learn more realistic parking behavior patterns in high-demand areas, thereby improving the prediction accuracy for these parking lot nodes. In contrast, the original adjacency matrix, with its uniform and unstructured weight distribution, results in concentrated errors in important parking lot areas during training, further hindering the model’s overall prediction performance.
IV-I Comparison of Computational Complexity
TABLE VI: Comparison of Time and Space Complexity Across Different Parking Network Scales
1800Nodes 3600Nodes 5400Nodes 7200Nodes
FLOPs MU/Byte FLOPs MU/Byte FLOPs MU/Byte FLOPs MU/Byte
STSGCN 83.27G 297.59M 198.54G 789.69M 277.8G 1517.25M 395.07G 2467.30M
PRGAT+Coarsening+STSGCN+TCN-AE 52.31G 350.16M 127.29G 955.78M 191.56G 1917.7M 274.69G 3082.93M
To gain a deeper understanding of the time and space complexity of our framework when training on large-scale urban parking networks, we conducted a quantification experiment on computational cost (FLOPs) and memory usage (MU). In the evaluation process, we selected STSGCN as the baseline and set the graph coarsening rate to 0.6. The specific computational complexity comparison results are shown in Table \@slowromancapvi@.
As shown in Table \@slowromancapvi@, in terms of time complexity, our framework reduces the computational cost of each forward and backward pass by an average of 33.2% when training on large-scale parking networks. This optimization is primarily attributed to the graph coarsening technique, which compresses the original parking network into a smaller graph while preserving the main structure and spatiotemporal dependencies of the original graph. As a result, the model only needs to compute fewer nodes and edges during each forward and backward pass. In contrast, traditional parking prediction models typically rely on full-graph computation, where each node aggregates information from all its neighboring nodes in each layer. As the graph size increases, this approach leads to a significant increase in computational cost, which severely impacts training efficiency.It is worth noting that, while we introduce the PRGAT, Graph Coarsening, and TCN-AE modules into our framework, which inevitably add extra computation, the impact of these additional computations on the overall training time is negligible and can be considered almost insignificant when compared to traditional models that perform spatiotemporal graph convolution directly on the original parking network scale.
Specifically, the computational cost of the PRGAT module is proportional to the graph size, but through the selective weighting mechanism, the model does not perform equal computations for all neighboring nodes. Instead, it focuses on the neighbors most relevant to the current node, thereby reducing redundant calculations and avoiding unnecessary overhead. The Graph Coarsening module’s computational cost mainly comes from node merging and adjacency matrix updates, but compared to the downstream spatiotemporal GCNs, its impact on the overall computational cost is minimal. Finally, the TCN-AE module’s computational cost does not increase exponentially with the graph size but remains relatively stable, meaning it will not become a computational bottleneck during training.
In terms of space complexity, our framework does not show significant improvement in memory usage when processing large-scale parking networks. In fact, it even leads to an average increase of 22% in memory consumption. This increase mainly comes from the following two factors: 1. Graph Coarsening, the increased memory consumption in this module is primarily due to the need to store multiple intermediate results. For example, it requires maintaining the mapping between the original and coarsened graphs, storing clustering labels and cluster centers during the clustering process, and storing intermediate results such as adjacency matrices and Laplacian matrices for each layer of multi-level graph coarsening. 2. TCN-AE, this module requires storing a large amount of intermediate activation values and hidden states during feature compression and reconstruction. Particularly when handling long time-series data, AE needs to store the hidden states and corresponding activation values for each time step, which results in significant memory consumption.
This experiment indicates that while our framework effectively reduces computational costs, it inevitably increases memory consumption due to the introduction of techniques such as graph coarsening. However, we consider this additional memory overhead acceptable, as it is necessary for optimizing computational performance and improving the model’s prediction accuracy. Current parking prediction tasks generally have low requirements for memory storage, especially when handling large-scale spatiotemporal data, where memory usage typically does not become a bottleneck. In contrast, time efficiency has a more significant impact on parking prediction performance, as increased computation time often significantly affects model deployment and real-time prediction capabilities. Therefore, we adopted a ”trading space for time” optimization strategy, leveraging moderate memory costs to achieve more efficient computational performance. This approach enables the framework to effectively address large-scale urban parking prediction tasks, enhancing training efficiency and prediction accuracy while maintaining high precision.